OpenStudy (anonymous):

graph the equation using slope and y intercept. x+4y=16

OpenStudy (texaschic101):

do you know how to find the slope and y intercept ?

OpenStudy (abmon98):

4y=x+16 y=1/4x+4 1/4 is the slope and 4 is the y-intercept

OpenStudy (texaschic101):

If you put your equation in y = mx + b form, the slope is in the m position and the y intercept is in the b position. So first put the equation in y = mx + b form

OpenStudy (texaschic101):

x + 4y = 16 4y = -x + 16 y = -1/4x + 4 slope = -1/4 y intercept is (0,4) to graph : start at y intercept (0,4)....slope is -1/4, so you come down 1 (because it is negative) and then to the right 4 (because it is positive), then down 1, then to the right 4. Now the x intercpet can be found by subbing in 0 for y. x + 4y = 16 x + 4(0) = 16 x = 16....x intercept is (16,0) ....keep following your slope and you will cross the x axis at (16,0)